OpenStudy (popgoesthweasal123):

Solve for d. 14d+34(d+16)=−5

Awolflover1 (awolflover1):

\[48d+544=−5\]

OpenStudy (danieldjpon3):

@Awolflover1

OpenStudy (danieldjpon3):

Y u have same picture as @AloneS ??? :O

OpenStudy (danieldjpon3):

... :O

Awolflover1 (awolflover1):

Subtract 544 on both sides \[48d+544−544=−5−544\]

OpenStudy (danieldjpon3):

wait wait wait @Awolflover1 answer my question pls

OpenStudy (danieldjpon3):

How and Why?!?

OpenStudy (danieldjpon3):

IM SO CONFUSED!!!

OpenStudy (danieldjpon3):

My brain going to explode :P

OpenStudy (danieldjpon3):

...

OpenStudy (danieldjpon3):

im scared

OpenStudy (danieldjpon3):

D;

Awolflover1 (awolflover1):

\[48d=−549\]

OpenStudy (popgoesthweasal123):

get out danild if your not going to answer my qestion

OpenStudy (danieldjpon3):

I will now

OpenStudy (danieldjpon3):

And you can't make me get out btw

OpenStudy (popgoesthweasal123):

thank you

OpenStudy (danieldjpon3):

:D

OpenStudy (popgoesthweasal123):

but i can ask you to

OpenStudy (danieldjpon3):

But I can say No... XD

OpenStudy (danieldjpon3):

lmao

Awolflover1 (awolflover1):

Divide both sides by 48. \[\frac{ 48d }{ 48}=\frac{ -549 }{ 48}\]

OpenStudy (danieldjpon3):

im done, im out. You creep me out @Awolflover1 and @AloneS

OpenStudy (danieldjpon3):

pce out guys

OpenStudy (popgoesthweasal123):

than leave

OpenStudy (danieldjpon3):

ok

Awolflover1 (awolflover1):

now d= ?

OpenStudy (popgoesthweasal123):

i have no idea

Awolflover1 (awolflover1):

Ok let's try again then \[14d+34(d+16)=−5\]

OpenStudy (popgoesthweasal123):

ok

Awolflover1 (awolflover1):

Simplify \[48d+544=−5\]

OpenStudy (popgoesthweasal123):

ok

Awolflover1 (awolflover1):

Subtract 544 from both sides. \[48d+544−544=−5−544\]

OpenStudy (popgoesthweasal123):

-549 and 592d

OpenStudy (popgoesthweasal123):

maybe

OpenStudy (nick63689):

you still need help?

Awolflover1 (awolflover1):

close :-) \[48d=−549\]

OpenStudy (popgoesthweasal123):

now what

Awolflover1 (awolflover1):

Divide both sides by 48 \[\frac{ 48d }{ 48}=\frac{ -549 }{ 48}\]

OpenStudy (popgoesthweasal123):

1 and 11

Awolflover1 (awolflover1):

1 and 11 ?

OpenStudy (popgoesthweasal123):

i dont know

OpenStudy (popgoesthweasal123):